# SDXL-Lightning: Progressive Adversarial Diffusion Distillation

### 21 Feb 2024 v1

@@ 1 Introduction
Conceptually, the generation involves a flow that gradually transports samples between the data and the noise distribution. The diffusion model learns to predict the gradient at any location of this flow. The generation is simply transporting samples from the noise distribution to the data distribution by following the predicted gradient through the flow. Because the flow is complex and curved, the generation must take a small step at a time. Formally, the flow is expressed as ordinary differential equations (ODEs) [67]. In practice, generating a high-quality data sample requires more than 50 inference steps.
@@
Our method combines the best of both worlds from progressive [54] and adversarial distillation [58]. Progressive distillation ensures that the distilled model follows the same ODE flow and has the same mode coverage as the original model. However, progressive distillation with mean squared error (MSE) loss produces blurry results under 8 inference steps and we provide theoretical analysis in our paper. To mitigate the issue, we use adversarial loss at every stage of the distillation to strike a balance between quality and mode coverage. Progressive distillation also brings an additional benefit, i.e., for multi-step sampling, our model predicts the next location on the ODE trajectory instead of jumping to the ODE endpoints every time by other distillation approaches [66, 58, 75]. This better preserves the original model behavior and facilitates better compatibility with LoRA modules [16] and control plugins [76, 10, 74].
@@
Our distillation method produces new state-of-the-art SDXL models that support one-step/few-step generation at 1024px resolution. We open-source our distilled models as SDXL-Lightning.
@@ 2.4 Adversarial Distillation
SDXL-Turbo [58] is the latest and the most popular open-source model using adversarial diffusion distillation. It follows prior works [56, 57] to use a pre-trained image encoder DINOv2 [41] as the discriminator backbone to accelerate training. However, this brings several limitations. First, using an off-the-shelf vision encoder means it must operate in the pixel space instead of the latent space, which significantly increases computation, memory consumption, and training time, making high-resolution distillation impractical. This is likely the reason SDXL-Turbo only supports up to 512px resolution. Second, an off-the-shelf vision encoder only works at $t=0$. The distilled model has to be trained to jump to ODE endpoints $x_{0}$, but since the quality for one-step inference is not good enough, random noises are added again for multi-step inference. This way of multi-step inference significantly alters the model behavior, making it less compatible with existing LoRA modules [16] and control plugins [76, 10, 74]. Third, off-the-shelf encoders may be hard to find for other datasets (anime, line arts, etc.) and modalities (video, audio, etc.). This reduces the generalizability of the distillation method. Lastly, the adversarial objective alone does not force the model to follow the same ODE flow, so mode coverage is not enforced.
@@ 2.5 Other Distillation Methods
Consistency Model (CM) [66, 65] also requires jumping to the ODE endpoints at every inference step. This causes large model behavior changes for multi-step sampling which reduces compatibility with LoRA modules and plugins. This method has been applied to SDXL [36, 37] but its generation quality is poor under 8 steps. Consistency Trajectory Model (CTM) [22] adds adversarial loss and supports jumping to arbitrary flow locations, but the adversarial training is applied post-distillation, instead of during the distillation, and the method has not been applied to large-scale text-to-image models.
@@ 3.1 Why Distillation with MSE Fails
The learned ODE flow is determined by the forward schedule [29, 31], the loss function [27], and the model capacity. Given finite training samples, the underlying data distribution is ambiguous. The maximum likelihood estimation (MLE) is a distribution that assigns even probability only to the observed samples and zero everywhere else. If the model has infinite capacity, it will learn a flow of this maximum likelihood estimation and overfit to always produce observed samples and generate no new data. In practice, diffusion models can generate new data because neural networks are not perfect learners.
When the model is used in multi-step generations, it is stacked and has a higher Lipschitz constant and more non-linearities to approximate a more complex distribution. But when the model is used in few-step generations, it no longer has the same amount of capacity to approximate well the same distribution. This is evidenced by diffusion models can have very sharp changes in results despite small changes in the initial noises [9], but the distilled models have much smoother latent traversal.
This explains why distillation with MSE loss produces blurry results. The student model simply does not have the capacity to match the teacher. We find that other distance metrics, such as L1 and perceptual loss [77, 27], also produce undesirable results. On the other hand, we find adversarial objectives to be effective in mitigating this issue.
@@ 3.2 Adversarial Objective
The condition on $x_{t}$ is important for preserving the ODE flow. This is because the teacher’s generation of $x_{t-ns}$ is deterministic from $x_{t}$. By providing the discriminator both $x_{t-ns}$ and $x_{t}$, the discriminator learns the underlying ODE flow and the student must also follow the same flow to fool the discriminator.
Our formulation is very similar to a prior work [72] except we use it for distillation instead of training from scratch. Note that this approach only preserves the ODE flow and ensures mode coverage when used in distillation.
@@ 3.6 Distillation Procedure
Then, we switch to using adversarial loss to distill the step count in this order: $32\rightarrow 8\rightarrow 4\rightarrow 2\rightarrow 1$. At each stage, we first train with the conditional objective as in Section 3.2 to preserve the ODE flow, and then train with the unconditional objective as in Section 3.4 to relax the mode coverage.
@@ 3.7.2 Train Discriminator at Multiple Timesteps
Specifically, we first draw $t*\leftarrow\{10,250,500,750\}$ with uniform weighting 1:1:1:1 and sample a new noise $\epsilon*\sim\mathcal{N}(0,\mathbf{I})$. Then we apply the noise before passing it through the discriminator:

### 02 Mar 2024 v3

@@ 1 Introduction
Conceptually, the generation involves a probability flow that gradually transports samples between the data and the noise probability distribution. The diffusion model learns to predict the gradient at any location of this flow. The generation is simply transporting samples from the noise distribution to the data distribution by following the predicted gradient through the flow. Because the flow is complex and curved, the generation must take a small step at a time. Formally, the flow can be expressed as an ordinary differential equation (ODE) [67]. In practice, generating a high-quality data sample requires more than 50 inference steps.
@@
Our method combines the best of both worlds from progressive [54] and adversarial distillation [58]. Progressive distillation ensures that the distilled model follows the same probability flow and has the same mode coverage as the original model. However, progressive distillation with mean squared error (MSE) loss produces blurry results under 8 inference steps and we provide theoretical analysis in our paper. To mitigate the issue, we use adversarial loss at every stage of the distillation to strike a balance between quality and mode coverage. Progressive distillation also brings an additional benefit, i.e., for multi-step sampling, our model predicts the next location on the ODE trajectory instead of jumping to the ODE trajectory endpoints every time by other distillation approaches [66, 58, 75]. This better preserves the original model behavior and facilitates better compatibility with LoRA modules [16] and control plugins [76, 10, 74].
@@
Our distillation method produces new state-of-the-art SDXL [44] models that support one-step/few-step generation at 1024px resolution. We open-source our distilled models as SDXL-Lightning.
@@ 2.4 Adversarial Distillation
SDXL-Turbo [58] is the latest and the most popular open-source model using adversarial diffusion distillation. It follows prior works [56, 57] to use a pre-trained image encoder DINOv2 [41] as the discriminator backbone to accelerate training. However, this brings several limitations. First, using an off-the-shelf vision encoder means it must operate in the pixel space instead of the latent space, which significantly increases computation, memory consumption, and training time, making high-resolution distillation impractical. This is likely the reason SDXL-Turbo only supports up to 512px resolution. Second, an off-the-shelf vision encoder only works at $t=0$. The distilled model has to be trained to jump to ODE trajectory endpoints $x_{0}$, but since the quality for one-step inference is not good enough, random noises are added again for multi-step inference. This way of multi-step inference significantly alters the model behavior, making it less compatible with existing LoRA modules [16] and control plugins [76, 10, 74]. Third, off-the-shelf encoders may be hard to find for other datasets (anime, line arts, etc.) and modalities (video, audio, etc.). This reduces the generalizability of the distillation method. Lastly, the adversarial objective alone does not force the model to follow the same probability flow, so mode coverage is not enforced.
@@ 2.5 Other Distillation Methods
Consistency Model (CM) [66, 65] also requires jumping to the ODE trajectory endpoints at every inference step. This causes large model behavior changes for multi-step sampling which reduces compatibility with LoRA modules and plugins. This method has been applied to SDXL [36, 37] but its generation quality is poor under 8 steps. Consistency Trajectory Model (CTM) [22] adds adversarial loss and supports jumping to arbitrary flow locations, but the adversarial training is applied post-distillation, instead of during the distillation, and the method has not been applied to large-scale text-to-image models.
@@ 3.1 Why Distillation with MSE Fails
The learned probability flow is determined by the dataset, the forward function [29, 31], the loss function [27], and the model capacity. Given finite training samples, the underlying data distribution is ambiguous. The maximum likelihood estimation (MLE) is a distribution that assigns even probability only to the observed samples and zero everywhere else. If the model has infinite capacity, it will learn a flow of this maximum likelihood estimation and overfit to always produce observed samples and generate no new data. In practice, diffusion models can generate new data because neural networks are not exact learners.
When the model is used in multi-step generations, it is stacked and has a higher Lipschitz constant and more non-linearities to approximate a more complex distribution. But when the model is used in few-step generations, it no longer has the same amount of capacity to approximate well the same distribution. This is evidenced by diffusion models can have very sharp changes in results despite small changes in the initial noises [9], but the distilled models have much smoother latent traversal. This explains why distillation with MSE loss produces blurry results. The student model simply does not have the capacity to match the teacher.
Additionally, neural network parameter optimization involves a complex landscape. Even models with the same capacity can hardly match output exactly since parameters can get stuck at different local minima.
We find that other distance metrics, e.g. L1 and perceptual loss [77, 27], also produce undesirable results. On the other hand, we find adversarial objectives to be effective in mitigating this issue.
@@ 3.2 Adversarial Objective
The condition on $x_{t}$ is important for preserving the probability flow. This is because the teacher’s generation of $x_{t-ns}$ is deterministic from $x_{t}$. By providing the discriminator both $x_{t-ns}$ and $x_{t}$, the discriminator learns the underlying probability flow and the student must also follow the same flow to fool the discriminator.
Our formulation is very similar to a prior work [72] except we use it for distillation instead of training from scratch. Note that this approach only preserves the probability flow and ensures mode coverage when used in distillation.
@@ 3.6 Distillation Procedure
Then, we switch to using adversarial loss to distill the step count in this order: $32\rightarrow 8\rightarrow 4\rightarrow 2\rightarrow 1$. At each stage, we first train with the conditional objective as in Section 3.2 to preserve the probability flow, and then train with the unconditional objective as in Section 3.4 to relax the mode coverage.
@@ 3.7.2 Train Discriminator at Multiple Timesteps
Specifically, we first draw $t*\leftarrow\{10,250,500,750\}$ with uniform weighting 1:1:1:1 and sample a new noise $\epsilon*\sim\mathcal{N}(0,\mathbf{I})$. Then we apply the noise before passing it through the conditional and unconditional discriminators:
@@ 3.7.3 Switch to $x_{0}$ Prediction
As discussed in Section 3.1, MSE loss cannot convert our model perfectly. The converted model generates blurry results, but this will be fixed by the adversarial objectives.